"Jim Hart is 5 years younger than Joan Chaput. Five years ago Jim was two-thirds as old as Joan. How old are they?"
here's what I tried to do:
x = joan x-5 = jim... jim 5yrs. ago = x-5-5 joan 5yrs. ago = x-5
x-5-5 = 2/3(x-5)
then I distributed after converting the fraction to a decimal
x-5-5 = .66x - 3.3
then i tried to get the x's together on one side by adding ten to both sides
x = .66x + 6.7
here's where i get stuck,i don't know what to do next... i don't even know if i laid it out right. please help, thanks!

Answer by scott8148(6628)

x-5-5 = 2/3(x-5) ___ this is correct , but you might want to eliminate the fraction by multiplying by 3
___ it makes for a "less messy" solution

3x - 15 - 15 = 2x - 10

adding 30 ___ 3x = 2x + 20

subtracting 2x ___ x = 20